OpenStudy (anonymous):

Which of the following tables shows the correct steps to transform x2 + 6x + 8 = 0 into the form (x - p)2 = q? [p and q are integers]

OpenStudy (anonymous):

Step 1 x2 + 6x + 8 - 1 = 0 - 1 Step 2 x2 + 6x + 7 = -1 Step 3 (x + 3)2 = -1 Step 1 x2 + 6x + 8 + 1 = 0 + 1 Step 2 x2 + 6x + 9 = 1 Step 3 (x + 3)2 = 1 Step 1 x2 + 6x + 8 - 2 = 0 - 2 Step 2 x2 + 6x + 6 = -2 Step 3 (x + 3)2 = -2 Step 1 x2 + 6x + 8 + 2 = 0 + 2 Step 2 x2 + 6x + 10 = 2 Step 3 (x + 3)2 = 2

OpenStudy (anonymous):

@amistre64 @ganeshie8 @beccaboo333

OpenStudy (amistre64):

well, how would you go about it?

OpenStudy (amistre64):

what is (x+3)^2?

OpenStudy (anonymous):

what is it factored?

OpenStudy (amistre64):

no, what is (x+3)^2 when we expand it into a trinomial? if FOIL sounds familar then work it

OpenStudy (anonymous):

I haven't learned foil

OpenStudy (amistre64):

have you learned to multiply binomials?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

all FOIL is is some memory device used to remember how to multiply binomials .... i find it to be utterly useless myself. (x+3)^2 = (x+3)(x+3) = ____ + _____ + _____

OpenStudy (anonymous):

so multiply (x+3)(x+3)?

OpenStudy (amistre64):

of course ....

OpenStudy (amistre64):

we are simply working backwards to see what step2 needs to look like.

OpenStudy (anonymous):

x^2+6x+9

OpenStudy (amistre64):

correct, now which options has that as the correct step2?

OpenStudy (anonymous):

i don't see one that does

OpenStudy (anonymous):

wait nvm its the second one

OpenStudy (amistre64):

:) yeah

OpenStudy (anonymous):

okay thank you , do you think you could try and help with another question?

OpenStudy (amistre64):

maybe one more

OpenStudy (anonymous):

Functions f(x) and g(x) are shown below: f(x) = x2 g(x) = x2 - 10x + 25 In which direction and by how many units should f(x) be shifted to obtain g(x)? Right by 5 units Left by 5 units Right by 10 units Left by 10 units

OpenStudy (anonymous):

I take that as a no?

OpenStudy (amistre64):

take what as a no?

OpenStudy (amistre64):

let x=0, what do we get?

OpenStudy (anonymous):

for which one

OpenStudy (amistre64):

both of them ... we can compare the results. unless you know of a different way to approach it.

OpenStudy (amistre64):

do you know how to find a vertex yet?

OpenStudy (anonymous):

the first one is 0 the second one is 15 and yes

OpenStudy (anonymous):

its either the first or second one

OpenStudy (amistre64):

lets determine the x value of the vertex for g then, that should help determine where they moved it to :)

OpenStudy (anonymous):

(5,0) is the vertex for g

OpenStudy (amistre64):

then they moved f(0,0) to g(5,0) which option shows this movement?

OpenStudy (anonymous):

b?

OpenStudy (amistre64):

 i think its moved to the right by 5 .... as long as my directions arent messsed up lol

OpenStudy (anonymous):

gotcha, lol thank you!!

OpenStudy (amistre64):

youre welcome ;)

OpenStudy (squirrels):

@mathmorelikeno